2. Simplify each of the following expressions :

(i) (3 + 3 )(2 + 2 ) (ii) (3 + 3 )(3 − 3 )


(iii) ( 5+ 2)2

(iv) ( 5− 2)( 5+ 2)

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) (3+ 3)(2+ 2) =3#2+3 2 +2 3 + 6 = 6+3 2+2 3+ 6 (ii) (3 + 3 )(3 − 3 ) = (3)2 − ( 3 )2 = 9−3=6 (iii)( 5 + 2 )2 = ( 5 )2 + 2 $ 5 $ 2+( 2)2 = 5+2 10+2 = 7+2 10 (iv) ( 5− 2)( 5+ 2) =( 5)2−( 2)2 Class 9 • Maths • NCERT Solutions =5−2=3